Please help me with this problem.
The perimeter of a rectangle is 84 feet. The legnth of the rectangle is 3 feet less than 2 times its width. What is the legnth of this rectangle?
This is what I can do
84=2(w-3)
2(w-3)= 2w-6
2(2w-6)=4w-12
84=4w-12= 4w=84+12
4w=96
4/4=4/96
w=24
this is as far as I can get. I know the answer is soppose to be 27 feet but I can not solve for this answer.

Answer by Paul(988)

2(x+2x-3)=84
6x-6=84
6x=90
x=15
2(15)-3=27
Hence the width is 15ft and the length is 27ft